OpenStudy (anonymous):

studying for my exam Please help! solve: xC3 (combination) xPx-3 (permutation)

OpenStudy (anonymous):

if i recall correctly there is a formula for each. try to equate them resepectively.

OpenStudy (anonymous):

I know the formula, I'm just having a hard time solving these cause of the x

OpenStudy (anonymous):

i know that its X!/(x-3)!3! and 3!=6 but I'm not sure what to do with the parenthesis…because i domnt think i can distribute the !

OpenStudy (anonymous):

i think the (x-3)! = (x-3)(x-2)(x-1)(x-0) then you can distribute. But I dont think you can solve this because xC3 (combination) xPx-3 (permutation) Neither of them is equal to anything.

OpenStudy (anonymous):

ohh ok! well i do know there is an answer because i have it, just having a hard time to solve it.